# Bi-directional Deformation for Parameterization of Neural Implicit Surfaces

### v1

@@ 3.2 Laplacian Regularization
Mitigating angle distortion is crucial in ensuring the quality of parameterization. One approach to reduce angle distortion involves minimizing the ratio of the two singular values of the Jacobian matrix associated with the parameterization (Floater & Hormann, 2005).
However, considering the representation of the surface $\mathcal{S}$, the parametric domain $\mathcal{D}$, and the parameterization $F_{\text{def}}$ as neural implicit functions in our approach, directly computing the Jacobian matrix is a non-trivial task. To circumvent this, we employ a Laplacian regularizer in our implementation to control angle distortion. Achieving the minimizer of the Laplacian of the map results in a harmonic map. For genus-0 surfaces, harmonic maps are also conformal, which preserves angles (Gu et al., 2004). Although for surfaces of higher genus, a harmonic map is not necessarily conformal, it still serves an effective tool for reducing angle distortion (Gu & Yau, 2003). Specifically, for a sample point $\mathbf{p}\in\mathcal{S}$ and its neighboring points $\mathbf{q}_{j}\in\mathcal{S}$, the forward deformation maps them to $\mathbf{p}^{\prime}\in\mathcal{D}$ and $\mathbf{q}_{j}^{\prime}\in\mathcal{D}$, respectively. We compute the Laplacian of the sample $\mathbf{p}$ as follows:
$\Delta(\mathbf{p})=\sum_{j=1}^{m}\omega_{j}(\mathbf{p}^{\prime}-\mathbf{q}^{\prime}_{j}).$ (5)
@@
$\omega_{j}=\exp\left(-\frac{\|\mathbf{p}-\mathbf{q}_{j}\|_{2}}{l}\right),$ (6)
@@
$\mathcal{L}_{\text{Lap}}=\sum_{\mathbf{p}\in\mathcal{S}}\|\Delta(\mathbf{p})-\Delta(\mathbf{p})_{\perp}\|_{2}.$ (8)
@@ 3.3 Appearance Decomposition
Radiance represents the directional emission of color, illustrating how the appearance of an object varies in response to changes in the observational viewpoint (Mildenhall et al., 2020). Drawing inspiration from previous works (Ma et al., 2022; Ye et al., 2023) and applying the assumptions of Lambertian and grayscale shading (Fan et al., 2018), we decompose the radiance field of the target surface into two components: a view-independent material field and a view-dependent shading field (see Figure 7(b)). This decomposition allows for independent editing of material and shading. More precisely, for a sample point $\mathbf{p}\in\mathcal{S}$, we employ two sub-networks, $F_{\text{mat}}$ and $F_{\text{shd}}$, to decompose the radiance $\mathbf{r}\in\mathbb{R}^{3}$ at point $\mathbf{p}$ as follows:
@@
Then we employ the SDF-based volume rendering (Yariv et al., 2021) to integrate the radiances along the ray to compute color $\mathbf{c}_{\text{pred}}=\sum_{i}\lambda_{i}\mathbf{r}_{i}$, where $\mathbf{r}_{i}$ is the radiance of a sample point $\mathbf{p}_{i}$ and $\lambda_{i}$ is the same color weight as used in Equation (4).
GT Normal $\mathcal{D}$ Normal $\mathcal{D}$ Normal $\mathcal{D}$
Figure 3: We offer the flexibility to choose from a range of parametric domains for managing area distortion. From left to right: the closer the domain $\mathcal{D}$ to the 3D surface $\mathcal{S}$, the lower the area distortion in the parameterization.
@@ 3.4 Choice of Parametric Domains
When mapping a 3D surface to a parametric domain, angle and area distortions are usually inevitable. We use the Laplacian regularizer to reduce angle distortion. Given the existence of infinite conformal maps (Jin et al., 2004), identifying those with minimal area distortion is highly desired. Area distortion is influenced by the similarity between the 3D surface and the parametric domain: the closer their alignment, the less the distortion. Thus, a simplistic, closely aligned parametric domain is ideal, balancing simplicity and alignment. Our algorithm accommodates various domains such as spheres, cubes, and polycubes, allowing flexibility in selection. It should be noted that our method supports more complex parametric domains, such as polyhedral complexes Yang et al. (2023), thanks to the representation capability of the SDF sub-network. In our current implementation, we use spheres, cubes and polycubes, mainly because of their simplicity. Figure 3 illustrates the impact on the choice of parametric domain to the parameterization quality.
@@ 3.5 Training Losses
In contrast to NeP (Ma et al., 2022), which necessitate prior information such as tracked mesh and UV mapping for supervision, our method is capable of reconstructing 3D geometry and computing the corresponding texture map directly from multi-view images. This process is supervised by the image loss
@@
where $\mathbf{c}_{\text{gt}}$ and $\mathbf{c}_{\text{pred}}$ are the ground-truth and predicted colors, respectively. To ensure the shape and appearance latent codes, $\mathbf{z}_{s}$ and $\mathbf{z}_{a}$, conform to Gaussian distributions, we incorporate a regularization term, expressed as $\mathcal{L}_{\text{code}}=\|\mathbf{z}_{s}\|_{2}+\|\mathbf{z}_{a}\|_{2}$ (Park et al., 2019). Additionally, since both the 3D surface $\mathcal{S}$ and the parametric domain $\mathcal{D}$ are represented as signed distance fields, we employ the Eikonal loss (Yariv et al., 2020; Gropp et al., 2020), defined as $\mathcal{L}_{\text{Eik}}=(\|\nabla s\|_{2}-1)^{2}$. Putting it all together, we define the total loss function as follows:
$\mathcal{L}=\mathcal{L}_{\text{rgb}}+\lambda_{1}\mathcal{L}_{\text{Eik}}+\lambda_{2}\mathcal{L}_{\text{cycle}}+\lambda_{3}\mathcal{L}_{\text{smooth}}+\lambda_{4}\mathcal{L}_{\text{Lap}}+\lambda_{5}\mathcal{L}_{\text{shd}}+\lambda_{6}\mathcal{L}_{\text{code}}.$ (13)
@@
GT Normal $\mathcal{D}$ $\mathcal{S}$Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
GT Normal $\mathcal{D}$ $\mathcal{S}$Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
(a) (b)
Figure 4: Parameterization results. (a) Human heads are co-parameterized to a sphere owing to the clear geometric resemblance. (b) Given the diverse geometry of cars, each is parameterized to a polycube domain. From left to right: the input image, the normal map of the reconstructed geometry $\mathcal{S}$, the parametric domain $\mathcal{D}$, and the reconstructed surface $\mathcal{S}$. See Appendix A.2 for additional results.
@@ 4 Experiments
Datasets. For our experiments, we used a customized dataset named FS-Syn, derived from the Facescape dataset (Yang et al., 2020). From Facescale, we selected 10 human head models and configured 30 fixed viewpoints and lighting conditions for each to obtain synthesized multi-view images. We also evaluated our method on several man-made objects from the DTU dataset (Jensen et al., 2014) and the OmniObject dataset (Wu et al., 2023).
Training details. All the sub-networks $F_{\text{def}}$, $F_{\text{inv-def}}$, $F_{\text{sdf}}$, $F_{\text{mat}}$, and $F_{\text{shd}}$ are MLPs. See Appendix A.1 for the detailed network architectures. We trained our network using the Adam (Kingma & Ba, 2014) optimizer. Once the parametric domain is chosen, we first trained $F_{\text{sdf}}$ for 2,000 iterations to obtain the implicit neural representation of the parametric domain. Subsequently, we fixed $F_{\text{sdf}}$ and trained other sub-networks for 2,000 epochs driven by the loss function as Equation (13).
Results. Our neural parameterization framework enables the construction of a bijective map between objects and multiple parametric domains such as spheres, cubes, and polycubes, as shown in Figure 4. Additionally, our method supports co-parameterization of multiple objects and facilitates the transfer of material and shading between them. Our neural parameterization algorithm is fully compatible with existing neural rendering pipelines. Given multi-view images as input, we can reconstruct the 3D geometry and create the texture maps simultaneously. As shown in Figure 4, we can obtain high-fidelity results without any prior information.
UV prior    NeP w/o prior   NeP w/ prior         Ours
Figure 5: Parametrization results of NeP (Ma et al., 2022)
Compared with NeP (Ma et al., 2022), which is the latest neural surface editing work using neural parameterization, our method eliminates the need for prior information from UV mapping. Unlike our algorithm, NeP learns a map between surfaces and the UV plane. However, it heavily relies on an existing UV mapping as initialization, and yields rather poor texture mapping results without the UV mapping as prior (see Figure 5). Moreover, the results from NeP still exhibit significant distortions even with UV prior. Our method can effectively reduce angle distortion by minimizing Laplacian loss and reduce area distortion by choosing an appropriate parametric domain, as demonstrated in Figure 6 (left).
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption GT w/o $F_{\text{shd}}$ w/ $F_{\text{shd}}$
Figure 6: Ablation studies. Left: Utilization of the Laplacian loss has proven effective in reducing angle distortion, especially in regions with high curvature. Right: For scenes with significantly varying illumination, it is crucial to decompose view-dependent shading from the radiance field.
Radiance decomposition. In contrast to an entangled radiance field in NeRF (Mildenhall et al., 2020) and VolSDF (Yariv et al., 2021), we decompose the radiance into view-independent material and view-dependent shading according to Equation 9. We illustrate the decomposition results in Figure 7. We would like to emphasize that it is essential to decompose the radiance according to the view-dependent correlations for scenes with varying illumination. As shown in Figure 6 (right), it fails to reconstruct the 3D geometry when we remove the $F_{\text{shd}}$, and the visual quality drops subsequently.
GT Material Shading GT Material Shading
@@
Texture editing. Both NeuMesh (Yang et al., 2022) and NeP (Ma et al., 2022) can edit the object’s texture at a pixel level. We compare the editing and rendering results with NeuMesh and NeP, as shown in Figure 8. NeuMesh requires a mask of the editing region. It also re-trains the network after each edit, which is time-consuming and cannot guarantee the stability of the training results. Our method outperforms NeuMesh in terms of the visual quality. NeP, which is also neural parameterization-based method, requires a UV prior for initialization. It often yields poor parameterization results without such prior. Even with the UV prior, its parameterization results exhibit large distortion, making the edit in the 2D domain challenging. Our method is free of prior, and can yield parameterization with low distortion, thereby facilitating pixel-level editing and material transferring without network re-training.
GTRefer to captionRefer to caption
{linebox} Texture Ours NeuMesh NePRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption Texture editing
Refer to captionRefer to caption Texture transfer
@@
Before After         Before After
@@ 5 Conclusion
Our method has a few limitations that require further improvement in the future. First, computational results on the FS-Syn dataset show that the reconstruction quality of our method is slightly worse than the standard SDF-based neural rendering algorithms, such as VolSDF (Yariv et al., 2021) and NeuS (Wang et al., 2021). For instance, the PSNRs of our results and VolSDF’s are 30.87 and 31.15, respectively. This decline in visual quality arises from the incorporation of additional modules, such as bi-directional deformation and radiance decomposition, resulting in the cessation of updates to the SDF network within our neural parameterization pipeline. Therefore, computing a parameterization without compromising the reconstruction quality is highly desired. Second, our radiance decomposition operates under the assumptions of Lambertian reflectance and grayscale shading, which might hinder its effectiveness when dealing with intricate materials or sophisticated shadings. Third, in our current implementation, we did not optimize the runtime performance. Therefore, training our model is time consuming, requiring 3 days for the 10 human heads of the FS-Syn dataset (Yang et al., 2020). Incorporating recent advancements such as PermutoSDF (Rosu & Behnke, 2023) and Strivec (Gao et al., 2023) has the potential to significantly reduce the training and inference time.
Appendix A Appendix
A.1 Implementation details

### v2

@@ 3.2 Laplacian Regularization
Mitigating angle distortion is crucial in ensuring the quality of parameterization. One approach to reduce angle distortion involves minimizing the ratio of the two singular values of the Jacobian matrix associated with the parameterization [6]. However, considering the representation of the surface $\mathcal{S}$, the parametric domain $\mathcal{D}$, and the parameterization $F_{\text{def}}$ as neural implicit functions in our approach, directly computing the Jacobian matrix is a non-trivial task. To circumvent this, we employ a Laplacian regularizer in our implementation to control angle distortion. Achieving the minimizer of the Laplacian of the map results in a harmonic map. For genus-0 surfaces, harmonic maps are also conformal, which preserves angles [13]. Although for surfaces of higher genus, a harmonic map is not necessarily conformal, it still serves an effective tool for reducing angle distortion [12]. Specifically, for a sample point $\mathbf{p}\in\mathcal{S}$ and its neighboring points $\mathbf{q}_{j}\in\mathcal{S}$, the forward deformation maps them to $\mathbf{p}^{\prime}\in\mathcal{D}$ and $\mathbf{q}_{j}^{\prime}\in\mathcal{D}$, respectively. We compute the Laplacian of the sample $\mathbf{p}$ as follows:
$\Delta(\mathbf{p})=\sum\nolimits_{j=1}^{m}\omega_{j}(\mathbf{p}^{\prime}-\mathbf{q}^{\prime}_{j}).$ (5)
@@
$\omega_{j}=\exp\left(-\|\mathbf{p}-\mathbf{q}_{j}\|_{2}/l\right),$ (6)
@@
$\mathcal{L}_{\text{Lap}}=\sum\nolimits_{\mathbf{p}\in\mathcal{S}}\|\Delta(\mathbf{p})-\Delta(\mathbf{p})_{\perp}\|_{2}.$ (8)
@@ 3.3 Appearance Decomposition
Radiance represents the directional emission of color, illustrating how the appearance of an object varies in response to changes in the observational viewpoint [29]. Drawing inspiration from previous works [27, 62] and applying the assumptions of Lambertian and grayscale shading [4], we decompose the radiance field of the target surface into two components: a view-independent material field and a view-dependent shading field (see Figure 9 (b)). This decomposition allows for independent editing of material and shading. More precisely, for a sample point $\mathbf{p}\in\mathcal{S}$, we employ two sub-networks, $F_{\text{mat}}$ and $F_{\text{shd}}$, to decompose the radiance $\mathbf{r}\in\mathbb{R}^{3}$ at point $\mathbf{p}$ as follows:
@@
Then we employ the SDF-based volume rendering [61] to integrate the radiances along the ray to compute color $\mathbf{c}_{\text{pred}}=\sum_{i}\lambda_{i}\mathbf{r}_{i}$, where $\mathbf{r}_{i}$ is the radiance of a sample point $\mathbf{p}_{i}$ and $\lambda_{i}$ is the same color weight as used in Equation (4).
Refer to captionRefer to captionRefer to captionRefer to caption (a) $\mathcal{S}$ (b) 2.505 (c) 6.094 (d) 1.633Refer to captionRefer to captionRefer to captionRefer to caption (a) $\mathcal{S}$ (b) 1.793 (c) 1.932 (d) 1.425
Figure 4: We offer the flexibility to choose from a range of parametric domains for managing area distortion. The closer the parametric domain $\mathcal{D}$((b-d)) to the (a) 3D surface $\mathcal{S}$, the lower the area distortion $\mathcal{E}_{\text{area}}$ (shown below the objects, with a detailed definition provided in Section 4) in the parameterization.
@@ 3.4 Choice of Parametric Domains
An ideal parameterization should be conformal and area-preserving to avoid any signal distortion. For general objects, the overall parameterization to 2D domains, such as disks or squares, leads to large distortion. Thus, the standard 2D texture mapping is to cut the surface into several patches, each of which can be parameterized with low distortion. However, this multi-chart approach inevitably produces seams, and the process of creating a good overall mapping also requires one to define suitable boundaries of the charts in the 2D domain. Therefore, unlike NeP [27], we map the neural implicit surfaces to 3D parametric domains and provide the selection of an appropriate domain shape to reduce area distortion. In addition, we use the Laplacian regularizer to reduce angle distortion. Given the existence of infinite conformal maps [18], identifying those with minimal area distortion is highly desired. Area distortion is influenced by the similarity between the 3D surface and the parametric domain: the closer their alignment, the less the distortion. Thus, a simplistic, closely aligned parametric domain is ideal, balancing simplicity and alignment. Our algorithm accommodates various domains such as spheres, cubes, and polycubes, allowing flexibility in selection. It should be noted that our method supports more complex parametric domains, such as polyhedral complexes [59], thanks to the representation capability of the SDF sub-network. In our current implementation, we use spheres, cubes and polycubes, mainly because of their simplicity. Figure 4 illustrates the impact on the choice of parametric domain to the parameterization quality.
@@ 3.5 Training Losses
In contrast to NeP [27], which necessitate prior information such as tracked mesh and UV mapping for supervision, our method is capable of reconstructing 3D geometry and computing the corresponding texture map directly from multi-view images. This process is supervised by the image loss
@@
where $\mathbf{c}_{\text{gt}}$ and $\mathbf{c}_{\text{pred}}$ are the ground-truth and predicted colors, respectively. To ensure the shape and appearance latent codes, $\mathbf{z}_{s}$ and $\mathbf{z}_{a}$, conform to Gaussian distributions, we incorporate a regularization term, expressed as $\mathcal{L}_{\text{code}}=\|\mathbf{z}_{s}\|_{2}+\|\mathbf{z}_{a}\|_{2}$ [33]. Additionally, since both the 3D surface $\mathcal{S}$ and the parametric domain $\mathcal{D}$ are represented as signed distance fields, we employ the Eikonal loss [60, 10], defined as $\mathcal{L}_{\text{Eik}}=(\|\nabla s\|_{2}-1)^{2}$. Putting it all together, we define the total loss function as follows:
$\begin{split}\mathcal{L}=\mathcal{L}_{\text{rgb}}+&\lambda_{1}\mathcal{L}_{\text{Eik}}+\lambda_{2}\mathcal{L}_{\text{cycle}}+\lambda_{3}\mathcal{L}_{\text{smooth}}\\ &+\lambda_{4}\mathcal{L}_{\text{Lap}}+\lambda_{5}\mathcal{L}_{\text{shd}}+\lambda_{6}\mathcal{L}_{\text{code}}.\end{split}$ (13)
@@ 4 Experiments
Datasets. For our experiments, we used a customized dataset named FS-Syn, derived from the Facescape dataset [58]. From Facescale, we selected 10 human head models and configured 30 fixed viewpoints and lighting conditions for each to obtain synthesized multi-view images. We also evaluated our method on several man-made objects from the DTU dataset [17] and the OmniObject dataset [52].
Training details and metric. All the sub-networks $F_{\text{def}}$, $F_{\text{inv-def}}$, $F_{\text{sdf}}$, $F_{\text{mat}}$, and $F_{\text{shd}}$ are MLPs. See supplementary material for other detailed network hyperparameters. We trained our network using the Adam [19] optimizer. Once the parametric domain is chosen, we first trained $F_{\text{sdf}}$ for 2,000 iterations to obtain the implicit neural representation of the parametric domain. Subsequently, we fixed $F_{\text{sdf}}$ and trained other sub-networks for 2,000 epochs driven by the loss function as Equation (13). The quality of the polycube map is measured by the angle distortion and area distortion [3, 15] $\mathcal{E}_{\text{angle}},\mathcal{E}_{\text{area}}$ as follows:
$\mathcal{E}_{\text{angle}}=\frac{\cot\alpha|a|^{2}+\cot\beta|b|^{2}+\cot\gamma|c|^{2}}{4\text{area}(\Delta_{i})},$ (14)
$\mathcal{E}_{\text{area}}=\frac{1}{2}\left(\frac{\text{area}(\Delta_{i})}{\text{area}(\hat{\Delta}_{i})}+\frac{\text{area}(\hat{\Delta}_{i})}{\text{area}(\Delta_{i})}\right),$ (15)
where $a,b,c$ represent the side lengths of the triangle $\Delta_{i}$ of the original surface. $\alpha,\beta,\gamma$ are angles of the corresponding triangle $\hat{\Delta}_{i}$ in parametric domain.In the case of conformal mapping, $\mathcal{E}_{\text{angle}}$ is equal to 1, and $\mathcal{E}_{\text{area}}$ is also equal to 1 for an isometric map.
$\mathcal{S}$ $\mathcal{D}$
Figure 5: Parameterization results of objects with non-zero genus topology and complex shape.
Results. Our neural parameterization framework enables the construction of a bijective map between objects and multiple parametric domains such as spheres, cubes, and polycubes, as shown in Figure 3 and Figure 5. Additionally, our method supports co-parameterization of multiple objects and facilitates the transfer of material and shading between them. Our neural parameterization algorithm is fully compatible with existing neural rendering pipelines. Given multi-view images as input, we can reconstruct the 3D geometry and create the texture maps simultaneously. As shown in Figure 3, we can obtain high-fidelity results without any prior information.
UV prior NeP w/o prior w/ prior       Ours
Figure 6: Parametrization results of NeP [27]
Compared with NeP [27], which is the latest neural surface editing work using neural parameterization, our method eliminates the need for prior information from UV mapping. Unlike our algorithm, NeP learns a map between surfaces and the UV plane. However, it heavily relies on an existing UV mapping as initialization, and yields rather poor texture mapping results without the UV mapping as prior (see Figure 6). Moreover, the results from NeP still exhibit significant distortions even with UV prior. Our method can effectively reduce angle distortion by minimizing Laplacian loss and reduce area distortion by choosing an appropriate parametric domain. As demonstrated in Figure 7, we calculate the angle distortion $\mathcal{E}_{\text{angle}}$ of parameterization as described in [3, 15]. We can observe that the Laplacian loss can effectively reduce angle distortion.
@@
Method $\mathcal{E}_{\text{angle}}$downarrow
w/o $\mathcal{L}_{\text{Lap}}$ 1.147
w/ $\mathcal{L}_{\text{Lap}}$ 1.092
Figure 7: Ablation study. Utilization of the Laplacian loss has proven effective in reducing angle distortion ($\mathcal{E}_{\text{angle}}$), especially in regions with high curvature.
GT w/o $F_{\text{shd}}$ w/ $F_{\text{shd}}$
Figure 8: Ablation study. For scenes with significantly varying illumination, it is crucial to decompose view-dependent shading from the radiance field to further enable separate editing.
Radiance decomposition. In contrast to an entangled radiance field in NeRF [29] and VolSDF [61], we decompose the radiance into view-independent material and view-dependent shading according to Equation 9. We illustrate the decomposition results in Figure 9. We would like to emphasize that it is essential to decompose the radiance according to the view-dependent correlations for scenes with varying illumination. As shown in Figure 8 (down), it fails to reconstruct the 3D geometry when we remove the $F_{\text{shd}}$, and the visual quality drops subsequently.
GT Material Shading
@@
Texture editing. Both NeuMesh [57] and NeP [27] can edit the object’s texture at a pixel level. We compare the editing and rendering results with NeuMesh and NeP, as shown in Figure 10. NeuMesh requires a mask of the editing region. It also re-trains the network after each edit, which is time-consuming and cannot guarantee the stability of the training results. Our method outperforms NeuMesh in terms of the visual quality. NeP, which is also neural parameterization-based method, requires a UV prior for initialization. It often yields poor parameterization results without such prior. Even with the UV prior, its parameterization results exhibit large distortion, making the edit in the 2D domain challenging. Our method is free of prior, and can yield parameterization with low distortion, thereby facilitating pixel-level editing and material transferring without network re-training.
@@
Before After
@@ 5 Conclusion
GT VolSDF Ours
Figure 12: Geometry reconstruction results of VolSDF [61] and ours. Our method is capable of editing through parameterization, accompanied by an acceptable compromise in reconstruction quality.
Our method has a few limitations that require further improvement in the future. First, computational results on the FS-Syn dataset show that the reconstruction quality of our method is slightly worse than the standard SDF-based neural rendering algorithms, such as VolSDF [61] and NeuS [47]. For example, as illustrated in Fig. 12 and the right table, We report PSNR and CD compared with VolSDF [61] on Fs-Syn dataset [58].
Method PSNR CD($10^{-3}$)
VolSDF 31.15 3.270
Ours 30.87 3.423
@@
This decline in visual quality arises from the incorporation of additional modules, such as bi-directional deformation and radiance decomposition, resulting in the cessation of updates to the SDF network within our neural parameterization pipeline. Therefore, computing a parameterization without compromising the reconstruction quality is highly desired. Second, our radiance decomposition operates under the assumptions of Lambertian reflectance and grayscale shading, which might hinder its effectiveness when dealing with intricate materials or sophisticated shadings. Third, in our current implementation, we did not optimize the runtime performance. Therefore, training our model is time consuming, requiring 3 days for the 10 human heads of the FS-Syn dataset [58]. Incorporating recent advancements such as PermutoSDF [38] and Strivec [7] has the potential to significantly reduce the training and inference time.
Supplementary Material
Appendix A Implementation Details